Status of the shipment of small & reusable components to France
Status in a nutshell
- Container not yet at slac, nor ordered.
- E.H. cleared of DIRC components.
- 621 cabinets half-empty; the main items still stored are: mechanical tools + hardware related to the SOB, the spare DCCs and various FEE test bench items (spare components, tools, cables, etc.).
- HV cables going from the EH to the detector to be cut.
- Main task remaining: front-end boards and crates in 620 (inventory, removal and packing).
- Box BBR-STORAGE-003 (Grey, plastic, ex-Nasa)
- Hardware DB tag: 18016587
- 59 HV boards (LLR)
- 1 HV crate (LLR tag #00309; LLR)
- 2 light generators (Saclay)
- Box BBR-STORAGE-004 (White, wood, ex-Lockeed Martin)
- Hardware DB tag: 18016591
- The Apple G4 FEE teststand computer (LAL)
- 3 VSC boards (LAL)
- 2 PATGEN boards (LAL)
- 17 spare DFBs (LAL)
- The Bit3 board and its associated black cable (LAL)
- 1 spare Wiener PS (BBR tag #20004846; LAL)
- The 6 remaining HV crates (LLR)
- Misc. HV spare parts (cables, resistors, crate keys, connectors; LLR)
- 3 boxes of spare PMTs (Saclay)
- 5 boxes of spare PMT bases (Saclay)
- The spare LED booster (Saclay)
Last modified: Tue Dec 9 16:43:17 PST 2008